OpenStudy (anonymous):

find the (absolute) maximum value of the function f(x)=5x^2+x^3-x^4

OpenStudy (anonymous):

@terenzreignz hey can u please help me i only need to understand this one

terenzreignz (terenzreignz):

You need to differentiate it first, can you do that?

OpenStudy (anonymous):

no i dont know any of this please

terenzreignz (terenzreignz):

You don't know how to take a derivative?

OpenStudy (anonymous):

f'(x)=-4x^3+3x^2+10x this it

terenzreignz (terenzreignz):

Oh okay, you had me worried there ^_^ Now, equate this to zero, and solve for x. \[\Large -4x^3 +3x^2 +10x =0\] You should get three values...

OpenStudy (anonymous):

because i only get 12

terenzreignz (terenzreignz):

Solving for x? That's intermediate Algebra, isn't it?

OpenStudy (anonymous):

no i mean -1.25

terenzreignz (terenzreignz):

Okay, how did you get that?

OpenStudy (anonymous):

factoring

terenzreignz (terenzreignz):

True, but there ARE two other solutions...

OpenStudy (anonymous):

2 and 0

terenzreignz (terenzreignz):

Okay, good. Now, if your function has any maxima, it has to be at a point where the derivative was zero... Now... compare them... plug in zero, -1.25, and 2 to your function, and see which one gives you the largest value...

OpenStudy (anonymous):

2=12 -1.25=3.417 0=0

terenzreignz (terenzreignz):

So which is biggest?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

So its absolute maximum is...? (It's the largest value the function attains)

terenzreignz (terenzreignz):

And it's not 2 = 12, silly... it's f(2) = 12

OpenStudy (anonymous):

ok so is it 2 or 12 because they both are my answers choices

terenzreignz (terenzreignz):

Which do you think? What's the highest value the function attains...

OpenStudy (anonymous):

12

terenzreignz (terenzreignz):

Well then ;)

OpenStudy (anonymous):

thank u so much :)